Grade 12th passAlgebra

x*e^-x +1=0. find no.of roots in {-1,-1/2}

John Doe
12 Years agoGrade 12th pass

1 Answer

Arun Kumar
ApprovedApproved Tutor Answer12 Years ago
Hello Student,

f(x)=x*e^(-x)+1
=>f(-1)=-e+1 <0
=>f(-1/2)=1/2e^(-1/2)+1 >0
because f(x) is monotonic
=>only one root

Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty